Hi need some help with this problem, little bit confused.
I am to determine if the relationship represented in the table is linear. If it is linear, how would I write an equation?
x | 10 | 15 | 20 | 25 | 30 | 35 |
y | 10 | 20 | 30 | 40 | 40 | 50 |
I believe this is not a linear but if not how would I know.
Some of the choice I am looking at if it is linear would any of these fit.
y = 15x + 10
y = 10x + 20
y = 20x + 10
Thanks a bunch :o)

Answer by ewatrrr(24785)


Hi
Ordered pairs: (10,10),(15,20),(20,30),(25,40),(30,40),(35,50)
(15,20)
(10,10) m = 10/5 = 2 and
y = 2x + b
10 = 2*10 + b
-10 = b |y = 2x -10
1) (For the 1st 4 ordered pairs y =2x-10 works)..however NOT for the last two
These ordered pairs do not have a linear relationship